A vacuum recorded in the condenser of a steam power plant is 73kPa.Find the absolute pressure in the condenser in Pa. The barometric reading is 98kPa
1
Expert's answer
2021-10-01T12:21:23-0400

vacuum recorded 73kPa


absolute pressure= (Palm) - (Vacuum Press)=

=(98-73) kPa

=25 kPa